OpenStudy (anonymous):

A cell phone plan offers a monthly rate of $25 for unlimited minutes plus $.05 per text message. Which algebraic expression below represents the total cost for this cell phone plan in terms of "x?" 25.05x 25x + 0.05 25 + 0.05x 24.95x

OpenStudy (anonymous):

if x represents text message number, then there is a fix cost = 25 plus a cost per messgae. So 25+0,05x

OpenStudy (anonymous):

how bout this one Layla has allotted $76 for gifts while on vacation. She plans to buy a souvenir for her parents for $16 and spend the rest on 4 T-shirts for her friends. If x represents the price, in dollars, of a T-shirt, which inequality correctly represents the money Layla will spend on souvenirs? 4x + 16 ≤ 76 16x + 4 ≤ 76 4x + 16 ≥ 76 16x + 4 ≥ 76

OpenStudy (anonymous):

4x + 16 ≤ 76

OpenStudy (anonymous):

last one lol. Identify the correct x-intercept and y-intercept of the equation 15x - 5y = 30. x-intercept = (-2, 0); y-intercept = (0, 6) x-intercept = (2, 0); y-intercept = (0, -6) x-intercept = (0, -6); y-intercept = (2, 0) x-intercept = (0, 6); y-intercept = (-2, 0)

OpenStudy (anonymous):

for x -intr make y=0 in equation and solve for x for y -intr make x=0 in equation and solve for y

OpenStudy (anonymous):

so its this right... x-intercept = (2, 0); y-intercept = (0, -6)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

thankss

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i got one more for u . The depth d(x), in feet, of a circular swimming pool at a point x feet from the edge is represented by the function below. If the depth varies from 1 foot to 4 feet, which inequality describes the domain of the function? d(x)=1+x/8 0 ≤ x ≤ 8 0 ≤ x ≤ 24 0 ≤ x ≤ 40 1 ≤ x ≤ 8

OpenStudy (anonymous):

min dept =1 solfind x that gives this value for d(x) sane for max value = 4

OpenStudy (anonymous):

i mean sove: 1=1+x/8 and 4=1+x/8

OpenStudy (anonymous):

solve*

OpenStudy (anonymous):

so its D

OpenStudy (anonymous):

0 ≤ x ≤ 24

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

how bout this Identify the correct slope and y-intercept of the equation 5x - 3y = -6 slope =5/3 ; y-intercept = (0, 2) slope = -5; y-intercept = (0, -6) slope = -3/5; y-intercept = (0, 2) slope = 5/3; y-intercept = (0, -6)

OpenStudy (anonymous):

i gona do all your homework?

OpenStudy (anonymous):

this is simmilar like befor, i explained you how to do this

OpenStudy (anonymous):

oh ight